# A Tale of Two DL Cities: When Library Tests Meet Compiler

### v1

@@ Abstract
Deep Learning (DL) compilers typically load a DL model and optimize it with intermediate representation. Existing DL compiler testing techniques mainly focus on model optimization stages, but rarely explore bug detection at the model loading stage. Effectively testing the model loading stage requires covering diverse usages of each DL operator from various DL libraries, which shares a common objective with DL library testing, indicating that the embedded knowledge in DL library tests is beneficial for testing the model loading stage of DL compilers. In this work, we propose Opera to extract such domain knowledge from the test inputs for DL libraries. Opera constructs diverse tests from the various test inputs for DL libraries (including the test inputs documented in DL libraries and those generated by recent fuzzers). In addition, it incorporates a diversity-based test prioritization strategy to migrate and execute those test inputs that are more likely to detect diverse bugs earlier. We considered three sources of tests in DL libraries for migration and used eight frontends from three DL compilers (e.g., TVM, TensorRT, and OpenVINO) for evaluation. Opera detected $170$ previously unknown bugs in total, $90$ of which have been confirmed/fixed by developers, demonstrating the effectiveness of such the migration-based idea. The test prioritization strategy in Opera improves testing efficiency with migrated tests by 11.9$\sim$47.4on average compared to general test prioritization strategies.
@@ I Introduction
Similar to traditional compilers [7, 8, 9, 10], DL compilers also contain bugs, which can compromise the reliability of both the compilers themselves and the models they produce. As reported [4], each stage of DL compilers contains a significant number of bugs, underscoring the need for comprehensive testing to ensure the quality of DL compilers. However, existing DL compiler testing techniques primarily focus on the two optimization stages ( 2 and 3) mentioned earlier, neglecting the model loading stage ( 1). Specifically, recent DL compiler testing techniques (such as HirGen [11], Tzer [12], and TVMFuzz [13]) mainly construct tests at either high-level or low-level IRs, bypassing the model loading stage.
@@
Motivated by the above observations, we conducted the first study to explore the potential of test migration. We develop a technique, called Opera (OPERator Adapter), to test DL compilers, especially the model loading stage. Specifically, we consider three sources of test inputs in DL libraries for migration, i.e., test inputs documented in DL libraries and test inputs generated using two recent fuzzers (DocTer [15] and DeepREL [16]). However, the direct adoption of most DL library tests for testing the model loading stage of DL compilers is not feasible due to differences in their input formats. Specifically, while DL compiler tests rely on pre-constructed DL models, DL library tests typically involve subtasks (e.g., gradient calculation and model design) in model construction, many of which cannot be represented in the form of DL models. To address this challenge, Opera first extracts instances of DL operators from each DL library test via code instrumentation and then packages each operator instance to a model (as a migrated test for DL compilers) with the aid of model generation templates for different DL libraries. Each operator instance represents a specific usage of the operator, encompassing an operator API and its corresponding parameter settings.
@@
Another challenge that Opera encounters is the significant cost considerations. This is primarily due to two factors: (1) the substantial volume of migrated tests originating from various migration sources (e.g., tests documented in DL libraries and those generated by automatic test generation tools for DL library testing), and (2) the frequent migration and execution of tests resulting from the frequent evolution of both DL libraries and DL compilers. To address this challenge, the component of test prioritization is designed in Opera, which prioritizes the migration and execution of tests that are more likely to uncover a diverse range of bugs in the model loading stage. After prioritization, more bugs can be detected within any given time budgets, thereby enhancing overall test efficiency. The test prioritization component takes into account the diversity of operator instances.
In this work, we applied Opera to test three popular DL compilers (i.e., TVM [1], TensorRT [2], and OpenVINO [3]). To balance evaluation cost and conclusion generality, for each compiler, we chose several popular DL libraries, i.e., PyTorch, Keras, and ONNX, from its supported frontends (responsible for model loading). In total, our study covered eight frontends across three DL compilers.
@@
In total, Opera detects $170$ previously unknown bugs by migrating the knowledge embedded in DL library tests, $90$ of which have been confirmed/fixed, while the state-of-the-art grammar-based DL compiler testing technique (NNSmith) detects only $18$ bugs within the same time budget. The results demonstrate the effectiveness of such a migration-based idea for testing the model loading stage of DL compilers. Furthermore, the test efficiency can be largely improved with the test prioritization component in Opera. On average across eight subjects, it improves Opera without special test prioritization by 13.1% and improves Opera incorporating the widely-used test prioritization strategies in general software testing by 11.9$\sim$47.4in terms of APFD (Average Percentage of Faults Detected) [17].
@@
  • We designed a test-migration-based technique (Opera), which integrates various migration sources (i.e., tests documented in DL libraries and those generated by recent fuzzers), along with diversity-based test prioritization.
  • We conducted an extensive study to systematically evaluate Opera across eight frontends from three DL compilers, leading to the efficient detection of $170$ previously unknown bugs.
@@ II Motivation
In the model loading process, DL compilers convert DL models into semantically equivalent high-level IRs by precisely converting each operator instance into a high-level IR node. Hence, testing for the model loading stage actually shares similar features to the tests for every single operator in DL libraries. Both of them aim to verify that the system under test properly manages each operator instance. This motivates the systematic study on the feasibility and effectiveness of migrating knowledge embedded in DL library tests to test the model loading stage.
@@ III Approach
Instead of manually designing test inputs or developing test generators for covering operators in various DL libraries, in this work, we conduct the first exploration to enhance the testing of the model loading stage in DL compilers by migrating test inputs in a nearby domain, i.e., DL library testing. Such a test-migration idea can help cover operators in various DL libraries in a lightweight manner. In particular, it can relieve the burden of understanding and handling constraints in various operators when designing test inputs or test generators, which has been addressed well in the existing test inputs of DL libraries.
With the test-migration idea, we propose a technique, called Opera. It consists of three main steps. Opera first collects and migrates the test inputs accumulated in the area of DL library testing through operator instance extraction (Section III-A). Due to the large number of migrated test inputs and the requirement of frequent test migration and execution caused by the evolution of DL libraries and compilers, Opera then prioritizes migrated test inputs based on their diversity in order to improve testing efficiency (Section III-B). That is, the test inputs that are more likely to detect diverse bugs can be executed earlier. Finally, Opera incorporates two test oracles to determine whether a migrated test input detects a bug in the model loading stage of a DL compiler (Section III-C).
@@ III-A2 Operator Instance Extraction
Although there are massive test inputs from the three sources, However, most DL library tests cannot be directly adopted to test DL compilers due to differences in their input format. Specifically, DL compilers take DL models as inputs, while the tests for DL libraries are often in the format of Python code, most of which lack a complete model structure, as shown in Figure 1(b). Hence, extracting DL models from DL library tests for DL compiler testing is non-trivial.
@@
Note that the primary functionality of the model loading stage lies in converting each operator in a DL model individually into an equivalent IR, known as single-operator equivalence conversion. As a result, Opera naturally creates single-operator models for testing. Moreover, single-operator models facilitate the follow-up bug de-duplication and localization processes.
@@ III-B1 Diversity of Operator Signatures
The diversity of an operator signature is characterized by its occurrence frequencies in test cases of both DL libraries and DL compilers. A high frequency of an operator’s signature in the tests of deep learning libraries but a low frequency in those of deep learning compilers suggests that while deep learning library developers place significant importance on this operator, deep learning compiler developers do not. It is more probable that such operator signature induces bugs in DL compilers. Hence, Opera assigns an operator $\textit{OP}_{i}$ a higher priority score if it occurs in the set of migrated tests for a DL library more frequently (denoted the occurrence times as $\textit{Num}\_\textit{DLL}\_\textit{OP}_{i}$) but occurs in the test suite equipped by the DL compiler more rarely (denoted the occurrence times as $\textit{Num}\_\textit{DLC}\_\textit{OP}_{i}$). With the intuitions, the priority score of $\textit{OP}_{i}$ is calculated by the ratio of $\textit{Num}\_\textit{DLL}\_\textit{OP}_{i}$ over $\textit{Num}\_\textit{DLC}\_\textit{OP}_{i}$.
@@ III-B2 Diversity of Parameter Settings
If the value space of a parameter is a limited set of concrete values, Opera treats each unique value as a unique subspace as each of them may represent a unique configuration of using this operator. If the value space of a parameter is a range of Integers, Opera pays more attention to some special values in DL [21, 22], i.e., -1 (e.g., it can be used to represent the automatically-calculated dimension size in tensor shape), 0 (e.g., it can mean that no padding operation is performed when assigned to the parameter padding), and 1 (e.g., it can be a boundary value for the parameter scale in keras.layers.Rescaling). Hence, Opera partitions them into five sets of subspaces, i.e., $(-\infty,-2],[-1],[0],[1]$, and $[2,\infty)$. Similarly, Opera partitions the value space of a range of Floating number for a parameter into three sets of subspaces, i.e., (,0),[0]0delimited-[]0($\infty$,0),[0]( - ∞ , 0 ) , [ 0 ], and $(0,\infty)$.
@@ III-C Test Oracles
Crash has been widely used in compiler testing [26, 14], which refers to an unexpected termination of compilation process. To avoid crashing a DL compiler due to invalid tests, Opera uses the DL library from which the tests are migrated to check their validity in advance. If the DL library crashes when executing a test, this test is considered invalid and thus discarded before testing the DL compiler. In addition, the crashes that produce error messages like “unsupported type” and “unsupported operator” are disregarded as the unsupported features are often not treated as bugs.
@@ IV-C Metrics
All bugs are detected on the latest versions of DL compilers, and thus we created a bug report for each unique bug and then submitted it to project maintainers. We counted the number of confirmed or fixed bugs by developers. Based on the feedback from developers, all of our submitted bugs are unique. This indicates the accuracy of our de-duplication method.
Besides, Opera includes a test prioritization component to improve the testing efficiency, and thus it is also important to investigate the testing efficiency of each technique. Here, we used two metrics to evaluate the effectiveness of each prioritization strategy. First, we measured the time spent on detecting each bug. The shorter the time is, the more effective the strategy is. Second, We adopted the widely-used metric of evaluating test prioritization, i.e., APFD (Average Percentage of Faults Detected) [17], to compare various test prioritization strategies following many existing studies [38, 35]. The calculation of APFD is shown in Formula 1:
@@ IV-E Process
For each studied frontend, we obtained a set of migrated sets from the three migration sources with the aid of Opera. We then tested each frontend with these migrated tests and recorded whether a test triggered a failure or not and the time spent on executing each test. For fair comparison, we applied NNSmith and COMET to generate DL models for testing each frontend for the same time budget as that used by Opera (including the time spent on test generation, migration, prioritization, and execution by Opera), respectively.
To answer RQ2, for each frontend, we constructed four variants of Opera by replacing its diversity-based test prioritization strategy with each of the four compared strategies, respectively. By applying each variant to test each frontend, we recorded the test result and execution time of each test, and calculated the APFD value. To reduce the influence of randomness and the running environment, we repeated our experiments for five times and calculated average results.
@@ V-A2 Root Causes of Detected Bugs
In the following, we present two bugs detected by the migrated tests. Figure 5 depicts an Incorrect Code Logic bug in the PyTorch frontend of TVM [39].
@@
In this bug, the Threshold operator from PyTorch was converted into the high-level IR of _op.nn.relu, which differs in its computation logic. Any non-zero value assigned to the parameter threshold or value of the Threshold operator will lead to wrong inference results. Indeed, a migrated test containing the operator instance torch.nn.Threshold(threshold=2, value=1), helps trigger this bug during compilation. A patch [40] was committed to fix this bug by correcting the conversion logic for the Threshold operator as shown in Figure 5.
@@ V-A3 Comparison with NNSmith and COMET
During the same testing time, the migrated tests by Opera detect $79$, $43$, $48$ bugs, while NNSmith detects $11$, $2$, $5$ bugs and COMET detects $6$, 00, $4$ bugs in TVM, TensorRT, OpenVINO, respectively. 15 of 18 bugs detected by NNSmith and 7 of 10 bugs detected by COMET are unique, that were not detected by Opera. All these unique bugs are in the optimization stages. Besides, all frontend bugs detected by NNSmish and COMET are also detected by Opera. Opera detected 164 unique bugs that were not detected by NNSmith and COMET, 156 of which are frontend bugs. The results demonstrate the superiority of Opera in testing the model loading stage and the complementarity between Opera and the existing DL compiler testing techniques such as NNSmith and COMET.
The major reason for the superiority of migrated tests by Opera over NNSmith and COMET in testing the model loading stage is that the latter two support only 75 and 72 operators while the former covers 477 operators in a lightweight manner. This implies a correlation between operator coverage and bug detection, i.e., higher operator coverage in testing is likely to detect more bugs in the model loading stage.
@@ VI Discussion
Threats to Validity. The threats to validity mainly lie in the subjects and metrics. To ensure generalizability of Opera, we considered eight frontends from the three popular DL compilers (i.e., TVM, TensorRT, and OpenVINO) for evaluation following the existing studies [11, 12]. There are also some metrics for evaluation test prioritization, such as APFDc [17] and RAUC-k [45]. In our work, we used the most widely-used APFD metric and showed the trend of the number of detected bugs with the testing process proceeding. Moreover, we also used the RAUC-k metric but put the results at our project homepage due to the space limit and consistent conclusions.

### v2

@@ Abstract
Deep Learning (DL) compilers typically load a DL model and optimize it with intermediate representation. Existing DL compiler testing techniques mainly focus on model optimization stages, but rarely explore bug detection at the model loading stage. Effectively testing the model loading stage requires covering diverse usages of each DL operator from various DL libraries, which shares a common objective with DL library testing, indicating that the embedded knowledge in DL library tests is beneficial for testing the model loading stage of DL compilers. With this idea, we propose Opera to migrate the knowledge embedded in DL library tests to test the model loading stage. Opera constructs diverse tests from various tests for DL libraries (including the tests documented in DL libraries and those generated by recent fuzzers). In total, we considered three sources of tests in DL libraries for migration. In addition, it incorporates a diversity-based test prioritization strategy to migrate and execute those tests that are more likely to detect diverse bugs earlier. We then used eight frontends from three DL compilers (e.g., TVM, TensorRT, and OpenVINO) for evaluation. Opera detected $170$ previously unknown bugs in total, $90$ of which have been confirmed/fixed by developers, demonstrating the effectiveness of such the migration-based idea. The test prioritization strategy in Opera improves testing efficiency with migrated tests by 11.9$\sim$47.4on average compared to general test prioritization strategies.
@@ I Introduction
Similar to traditional compilers [7, 8, 9, 10, 11, 12, 13], DL compilers also contain bugs, which can compromise the reliability of both the compilers themselves and the models they produce. As reported [4], each stage of DL compilers contains a significant number of bugs, underscoring the need for comprehensive testing to ensure the quality of DL compilers. However, existing DL compiler testing techniques primarily focus on the two optimization stages ( 2 and 3) mentioned earlier, neglecting the model loading stage ( 1). Specifically, recent DL compiler testing techniques (such as HirGen [14], Tzer [15], and TVMFuzz [16]) mainly construct tests at either high-level or low-level IRs, bypassing the model loading stage.
@@
However, due to the separate development of communities for DL compiler testing and DL library testing, there has been no systematic study to investigate the feasibility of migrating the knowledge embedded in DL library tests for testing the model loading stage. Hence, we performed the first exploration on the potential of the migration-based idea. Specifically, we design a migration-based technique, called Opera (OPERator Adapter), to test DL compilers (especially the model loading stage) by considering three sources of tests in DL libraries for migration, i.e., tests documented in DL libraries and tests generated by two recent fuzzers (DocTer [18] and DeepREL [19]).
@@
In fact, the direct adoption of most DL library tests for testing the model loading stage of DL compilers is not feasible due to differences in their input formats. Specifically, while DL compiler tests rely on pre-constructed DL models, DL library tests typically involve subtasks (e.g., gradient calculation and model design) in model construction, many of which cannot be represented in the form of DL models. To address this challenge, Opera first extracts instances of DL operators from each DL library test via code instrumentation and then packages each operator instance to a model (as a migrated test for DL compilers) with the aid of model generation templates for different DL libraries. Each operator instance represents a specific usage of the operator, encompassing an operator API and its corresponding parameter settings.
@@
Another challenge that Opera encounters is the significant cost consideration. This is primarily due to two factors: (1) the substantial volume of migrated tests originating from various migration sources, and (2) the frequent migration and execution of tests resulting from the frequent evolution of both DL libraries and DL compilers. To address this challenge, the component of test prioritization is designed in Opera, which prioritizes the migration and execution of tests that are more likely to uncover a diverse range of bugs in the model loading stage. After prioritization, more bugs can be detected within any given time budgets, thereby enhancing the overall test efficiency. The test prioritization component takes into account the diversity of operator instances.
@@
In this work, we applied Opera to test three popular DL compilers (i.e., TVM [1], TensorRT [2], and OpenVINO [3]). To balance evaluation cost and conclusion generality, for each compiler, we chose several popular DL libraries, i.e., PyTorch, Keras, and ONNX, from its supported frontends (responsible for model loading). In total, our study covered eight frontends across three DL compilers. In total, Opera detects $170$ previously unknown bugs by migrating the knowledge embedded in DL library tests, $90$ of which have been confirmed/fixed, while the state-of-the-art grammar-based DL compiler testing technique (NNSmith) detects only $18$ bugs within the same time budget. The results demonstrate the effectiveness of such a migration-based idea for testing the model loading stage of DL compilers. Furthermore, the test efficiency can be largely improved with the test prioritization component in Opera. On average across eight subjects, it improves Opera without special test prioritization by 13.1% and improves Opera incorporating the widely-used test prioritization strategies in general software testing by 11.9$\sim$47.4in terms of APFD (Average Percentage of Faults Detected) [20].
@@
  • We designed a migration-based technique (Opera), which integrates various migration sources (i.e., tests documented in DL libraries and those generated by recent fuzzers), along with diversity-based test prioritization.
  • We conducted an extensive study to evaluate Opera across eight frontends from three DL compilers, leading to the efficient detection of $170$ previously unknown bugs.
@@ II Motivation
In the model loading process, DL compilers need the ability to handle various usages of these operators, including all the combinations of parameters, for correct transformation from DL operators to high-level IR. Hence, the model loading stage of DL compilers actually shares a similar test objective with the tests for these operators in DL libraries. This motivates the idea of migrating knowledge embedded in DL library tests to test the model loading stage.
@@ III Approach
With the migration-based idea, we propose a technique, called Opera. The workflow of Opera is shown in Figure 3, which contains two main components: test migration and test prioritization. Specifically, Opera first creates tests for the model loading stage by migrating knowledge embedded in DL library tests via operator instance extraction (Section III-A). Due to the large number of migrated tests and the requirement of frequent test migration and execution caused by the evolution of DL libraries and compilers, Opera then prioritizes migrated tests based on their diversity in order to improve testing efficiency (Section III-B). Finally, Opera incorporates two test oracles to determine whether a migrated test detects a bug in the model loading stage of a DL compiler (Section III-C).
@@ III-A2 Operator Instance Extraction
Although there are massive tests from the three sources, However, most DL library tests cannot be directly adopted to test DL compilers due to differences in their input format. Specifically, DL compilers take DL models as inputs, while the tests for DL libraries are often in the format of Python code, most of which lack a complete model structure, as shown in Figure 1(b). Hence, extracting DL models from DL library tests for DL compiler testing is non-trivial.
@@
Note that the primary functionality of the model loading stage lies in converting each operator in a DL model individually into an equivalent IR, known as single-operator equivalence conversion. As a result, Opera naturally creates single-operator models for testing. Also, single-operator models facilitate follow-up bug de-duplication and localization.
@@ III-B1 Diversity of Operator Signatures
The tests with different operator signatures mean that they have diverse semantics. However, how to determine the order of the tests with different operator signatures is still a challenge. In Opera, we address it according to the following intuitions: (1) the number of tests with an operator signature is large in DL library testing, indicating that this operator receives more attention potentially due to its complicated implementation logic, more corner cases in it, etc; (2) the number of tests with an operator signature is small in the test suite equipped by the DL compiler (i.e., the test suite specific to the model loading stage), indicating that DL compiler developers still pay little attention to testing the transformation of this operator. Hence, Opera assigns an operator $\textit{OP}_{i}$ a higher priority score if it occurs in the set of migrated tests for a DL library more frequently (denoted the occurrence times as $\textit{Num}\_\textit{DLL}\_\textit{OP}_{i}$) but occurs in the test suite equipped by the DL compiler more rarely (denoted the occurrence times as $\textit{Num}\_\textit{DLC}\_\textit{OP}_{i}$). With the intuitions, the priority score of $\textit{OP}_{i}$ is calculated by the ratio of $\textit{Num}\_\textit{DLL}\_\textit{OP}_{i}$ over $\textit{Num}\_\textit{DLC}\_\textit{OP}_{i}$.
@@ III-B2 Diversity of Parameter Settings
If the value space of a parameter is a limited set of concrete values, Opera treats each unique value as a unique subspace as each of them may represent a unique configuration of using this operator. If the value space of a parameter is a range of Integers, Opera pays more attention to some special values in DL [24, 25, 26], i.e., -1 (e.g., it can be used to represent the automatically-calculated dimension size in tensor shape), 0 (e.g., it can mean that no padding operation is performed when assigned to the parameter padding), and 1 (e.g., it can be a boundary value for the parameter scale in keras.layers.Rescaling). Hence, Opera partitions them into five sets of subspaces, i.e., $(-\infty,-2],[-1],[0],[1]$, and $[2,\infty)$. Similarly, Opera partitions the value space of a range of Floating number for a parameter into three sets of subspaces, i.e., (,0),[0]0delimited-[]0($\infty$,0),[0]( - ∞ , 0 ) , [ 0 ], and $(0,\infty)$.
@@ III-C Test Oracles
Crash has been widely used in compiler testing [30, 17], which refers to an unexpected termination of the compilation process. To avoid crashing a DL compiler due to invalid tests, Opera uses the DL library from which the tests are migrated to check their validity in advance. If the DL library crashes when executing a test, this test is considered invalid and thus discarded before testing the DL compiler. In addition, the crashes that produce error messages like “unsupported type” and “unsupported operator” are disregarded as the unsupported features are often not treated as bugs.
@@ IV-C Metrics
All bugs are detected on the latest versions of DL compilers, and thus we created a bug report for each unique bug and then submitted it to project maintainers. We counted the number of confirmed or fixed bugs by developers. Based on the feedback from developers, all of our submitted bugs that have been confirmed are unique. This indicates the accuracy of our de-duplication method.
Besides, Opera includes a test prioritization component to improve the testing efficiency, and thus it is also important to investigate the testing efficiency of each technique. Here, we used two metrics to evaluate the effectiveness of each prioritization strategy. First, we measured the time spent on detecting each bug. The shorter the time is, the more effective the strategy is. Second, We adopted the widely-used metric of evaluating test prioritization, i.e., APFD (Average Percentage of Faults Detected) [20], to compare various test prioritization strategies following many existing studies [41, 38, 42, 43]. The calculation of APFD is shown in Formula 1:
@@ IV-E Process
For each studied frontend, we obtained a set of migrated sets from the three migration sources with the aid of Opera. We then tested each frontend with these migrated tests and recorded whether a test triggered a failure or not and the time spent on each test. For fair comparison, we applied NNSmith and COMET to generate DL models for testing each frontend for the same time budget as that used by Opera (including the time spent on test generation, migration, prioritization, and execution by Opera), respectively.
To answer RQ2, for each frontend, we constructed four variants of Opera by replacing its diversity-based test prioritization strategy with each of the four compared strategies, respectively. By applying each variant to test each frontend, we recorded the test result and testing time of each test, and calculated the APFD value. To reduce the influence of randomness and the running environment, we repeated our experiments for five times and calculated average results.
@@ V-A2 Root Causes of Detected Bugs
Next, we present two bugs detected by the migrated tests. Figure 5 depicts an Incorrect Code Logic bug in the PyTorch frontend of TVM [44].
@@
In this bug, the Threshold operator from PyTorch was converted into the high-level IR of _op.nn.relu, which differs in its computation logic. Any non-zero value assigned to the parameter threshold or value of the Threshold operator will lead to wrong inference results. Indeed, a migrated test containing the operator instance torch.nn.Threshold(threshold=2, value=1), helps trigger this bug during compilation. A patch [45] was committed to fix it by correcting the conversion logic for the Threshold operator as shown in Figure 5.
@@ V-A3 Comparison with NNSmith and COMET
During the same testing time, the migrated tests by Opera detect $79$, $43$, $48$ bugs, while NNSmith detects $11$, $2$, $5$ bugs and COMET detects $6$, 00, $4$ bugs in TVM, TensorRT, OpenVINO, respectively. 15 of 18 bugs detected by NNSmith and 7 of 10 bugs detected by COMET are unique, which were not detected by Opera. All these unique bugs are in the optimization stages. Besides, all frontend bugs detected by NNSmish and COMET are also detected by Opera. Opera detected 164 unique bugs that were not detected by NNSmith and COMET, 156 of which are frontend bugs. The results demonstrate the superiority of Opera in testing the model loading stage and the complementarity between Opera and the existing DL compiler testing techniques (NNSmith and COMET). The major reason for the superiority of migrated tests by Opera over NNSmith and COMET in testing the model loading stage is that the latter two support only 75 and 72 operators while the former covers 477 operators in a lightweight manner. This implies a correlation between operator coverage and bug detection, i.e., higher operator coverage in testing is likely to detect more bugs in the model loading stage.
@@ VI Discussion
Comparing with direct model generation. Although designing a technique to generate single-operator models directly based on existing test generation fuzzers (e.g., DocTer and DeepREL) guided by the two diversity metrics (presented in Section III-B) can be efficient, it has little influence on the overall DL compiler testing. This is because the most time-consuming step for Opera is test execution (including model compilation) rather than test generation, which accounts for over 95% of the total time. Additionally, separating the test prioritization step in Opera helps optimize the execution of tests from multiple sources (including human-written tests and the tests generated by different fuzzers), indicating a global optimization strategy. In contrast, the test prioritization conducted by a fuzzer at each iteration is a local optimization strategy and it can not migrate tests from human-written tests in DL libraries, which actually detected the most DL compiler bugs (presented in Section V-A4), for testing DL compilers. Hence, we proposed a migration-based testing technique (i.e., Opera) rather than designing a model generator directly, considering the generalizability and effectiveness.
@@
Threats to Validity. The threats to validity mainly lie in the subjects and metrics. To ensure generalizability of Opera, we considered eight frontends from the three popular DL compilers (i.e., TVM, TensorRT, and OpenVINO) for evaluation following the existing studies [14, 15]. There are also some metrics for evaluation test prioritization, such as APFDc [20] and RAUC-k [50, 51]. In our work, we used the most widely-used APFD metric and showed the trend of the number of detected bugs with the testing process proceeding. Moreover, we also used the RAUC-k metric but put the results at our project homepage due to the space limit and consistent conclusions.